# Theoretical study on the crystal structure of a bilayer nickel-oxychloride Sr3Ni2O5Cl2 and analysis on the occurrence of possible unconventional superconductivity

### v1

@@ I Introduction
Discovery of superconductivity with a maximum $T_{c}$ of above 80 K in a bilayer Ruddlesden-Popper nickelate La3Ni2O7 (Fig. 1(a)) under pressure has sparked enormous interest in the field of condensed matter physics [1]. Although under pressure, $T_{c}$ exceeding the liquid nitrogen boiling temperature, and also the origin of superconductivity likely being unconventional [2, 3, 4], have led to a huge amount of experimental and theoretical studies after the discovery [1, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 3, 4, 94, 95, 2, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136]. In fact, the possibility of superconductivity in this material was discussed theoretically even before its experimental discovery by some of the previous authors [137], but one viewpoint that was missing at that time was the precise crystal structure, namely, although a tetragonal symmetry was simply assumed in Ref. 137, the actual structural symmetry is orthorhombic at ambient pressure. There are now accumulating studies suggesting that tetragonal symmetry of the crystal structure may be important for the occurrence of superconductivity in La3Ni2O7 [38, 55, 54], which is accomplished by applying pressure. One speculation regarding the reason for this can be that orthorhombicity favors density waves, which suppresses superconductivity. Interestingly, also for the trilayer compound La4Ni3O10, the superconducting phase appears by applying a sufficient pressure and the tetragonal structure is realized [36, 52, 53]. Then, a natural motivation arises for searching materials that have tetragonal symmetry at ambient pressure, and also have electronic properties similar to that of La3Ni2O7 at high pressures.
@@
In this study, we theoretically investigate the electronic structure of Sr3Ni2O5Cl2 and its structural stability. Our phonon calculation shows that this compound with the $I4/mmm$ tetragonal structure is dynamically stable, i.e., does not have imaginary-frequency phonon modes, even at ambient pressure. The Ni-$d_{3z^{2}-r^{2}}$ orbital energy is pushed down by the crystal field where a half of apical O2- ions is replaced with a Cl- ion. By this effect, the Ni-$d_{3z^{2}-r^{2}}$ orbital becomes rather closer to the half-filling in Sr3Ni2O5Cl2 than La3Ni2O7 whereas the stoichiometric Ni-$3d$ band filling are decreased in Sr3Ni2O5Cl2 ($d^{7}$) compared with La3Ni2O7 ($d^{7.5}$). We also investigate superconductivity in this compound with the multi-orbital Hubbard model constructed from our first-principles calculation, using the fluctuation exchange (FLEX) approximation combined with the linearized Eliashberg equation. While orbital hybridization between the $t_{2g}$ and $e_{g}$ orbitals, which is enhanced in this characteristic crystal field, is detrimental for superconductivity, a nearly half-filled Ni-$d_{3z^{2}-r^{2}}$ orbital enhances superconductivity. In total, we find that Sr3Ni2O5Cl2 is a promising new candidate of bilayer-nickelate superconductors, which can possess even higher $T_{c}$ than pressurized La3Ni2O7, at ambient pressure. We also check the formation enthalpy for synthesis of this compound. Our calculation shows that the high-pressure synthesis can be a good way to experimentally produce Sr3Ni2O5Cl2.
Figure 1: (a) Crystal structure of La3Ni2O7 at high pressure. (b) Crystal structure and (c) phonon band dispersion for Sr3Ni2O5Cl2. Crystal structures in this paper were depicted using the VESTA software [140]. Phonon dispersions in this paper were depicted in the Brillouin zone of the conventional tetragonal lattice.
@@ II Methods
For first-principles calculation, we used the PBEsol exchange-correlation functional [141] and the projector augmented wave method [142] as implemented in Vienna ab initio Simulation Package (VASP) [143, 144, 145, 146]. Core-electron states in PAW potentials were [Kr]$4d^{10}$, [Ar]$3d^{10}$, [Ar], [Ne], and [He] for La, Sr, Ni, Cl, and O, respectively. We used a plane-wave cutoff energy of 600 eV for Kohn-Sham orbitals without including the spin-orbit coupling for simplicity. A $12\times 12\times 12$ ${\bm{k}}$-mesh was used for self-consistent-field calculations.
We performed structural optimization until the Hellmann-Feynman force becomes less than 0.01 eV Å-1 for each atom. To check the stability of the tetragonal ($I4/mmm$) phase, we calculated the phonon dispersion for several materials. For this purpose, we used the finite displacement method as implemented in the Phonopy [147] software in combination with VASP. We used a $4\times 4\times 1$ ${\bm{q}}$-mesh for a conventional tetragonal unit cell. For a $4\times 4\times 1$ supercell used for finite-displacement calculations, we used a $3\times 3\times 2$ ${\bm{k}}$-mesh.
@@ III.1 Electronic and phonon band structures for Sr3Ni2O5Cl2
As described in Introduction, a key idea in this study is to employ a mixed-anion strategy [160] for stabilizing the tetragonal $I4/mmm$ phase without applying the pressure. In fact, for Sr3Ni2O5Cl2, while the NiO2 plane is somewhat buckled; the equatorial Ni-O-Ni angle is 162 as shown in Fig. 1(b), we found that the tetragonal structure is dynamically stable as shown by our phonon calculation (Fig. 1(c)) where no imaginary-frequency modes appear without applying the pressure. This is in sharp contrast to the case of La3Ni2O7, where the tetragonal structure is unstable at ambient pressure and high pressure is required to stabilize it [38, 55, 54]. Our phonon calculations for La3Ni2O7 also support this feature as shown in Fig. 6, where the imaginary-frequency modes appear for 0 GPa while they are absent for 20 GPa. Although we also checked the stability of the tetragonal structure for Sr3Ni2O4Cl3, where all the apical oxygens were replaced with chlorines, we found that the tetragonal structure was unstable for this case as shown in Fig. 7.
@@
It is noteworthy that the mixed-anion strategy drastically changes the crystal field of a nickel ion. Figure 2 presents the electronic band structures and partial density of states (pDOS) for La3Ni2O7 under pressure of 20 GPa and Sr3Ni2O5Cl2 without applying the pressure shown with an orbital weight of each Ni-$d$ orbital. We summarize important differences between two materials as follows.
@@
The average valence number of the nickel atom is different: Ni2.5+ ($d^{7.5}$) for La3Ni2O7 and Ni3+ ($d^{7}$) for Sr3Ni2O5Cl2. Thus, the number of $d$ electrons for Sr3Ni2O5Cl2 is less than that for La3Ni2O7. Nevertheless, the Ni-$d_{3z^{2}-r^{2}}$ orbital is rather closer to the half-filling in Sr3Ni2O5Cl2 as shown in Figs. 2(b)(h). This is because of the characteristic crystal field for the nickel atom in Sr3Ni2O5Cl2, where a half of apical O2- ions is replaced with a Cl- ion, which pushes down the orbital energy of the Ni-$d_{3z^{2}-r^{2}}$ orbital. As a result, the number of $d_{x^{2}-y^{2}}$ electrons is decreased in Sr3Ni2O5Cl2. On the other hand, Ni-$d_{xy/yz/zx}$ orbitals, which we hereafter call them the $t_{2g}$ orbitals while it is not a rigorous term for this symmetry, raise their energy levels relative to the Ni-$d_{3z^{2}-r^{2}}$ orbital. This is because the present crystal field strongly pushes down the Ni-$d_{3z^{2}-r^{2}}$ orbital level as we have stated above. Consequently, the strong DOS peak of the $t_{2g}$ orbitals lies just below the Fermi energy. In addition, the crystal field is strongly asymmetric with respect to the $xy$ plane (i.e., $+z$ and $-z$ are strongly inequivalent), which allows a sizable hybridization among $e_{g}$ ($d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$) and $t_{2g}$ orbitals. In the next section, we will discuss how these characteristics can affect superconductivity in this system.
@@ III.2 Model construction
In this and the next sections, we compared Sr3Ni2O5Cl2 without applying the pressure and La3Ni2O7 under pressure of 20 GPa, both of which have the $I4/mmm$ tetragonal structure in our calculation. We constructed a ten-orbital model consisting of five $3d$ orbitals centered at two Ni sites per unit cell. Figure 3(a) shows superposed band structures given by first-principles calculation and Wannier interpolation, where precise fitting around the Fermi level was achieved. Important parameter values extracted by Wannierization are given in Table 1, compared with those of four-orbital models consisting only of $e_{g}$ orbitals. The smaller differences in extracted parameters between the ten- and four-orbital models for La3Ni2O7 originate from the fact that the $e_{g}$ and $t_{2g}$ bands are well separated compared with Sr3Ni2O5Cl2, where the $t_{2g}$ and $e_{g}$ bands are rather entangled, as shown in Fig. 2. We found a relatively large hopping amplitude between $t_{2g}$ and $e_{g}$ orbitals: e.g., the nearest-neighbor intralayer hopping along the $x$ direction between the $d_{x^{2}-y^{2}}$ and $d_{xz}$ orbitals and that between the $d_{3z^{2}-r^{2}}$ and $d_{xz}$ orbitals are $-0.296$ and $0.199$ eV, respectively. This is due to the strongly asymmetric crystal field for a NiO5Cl octahedron as mentioned in the previous section. It is also noteworthy that the relatively strong buckling of NiO2 plane in Sr3Ni2O5Cl2 seems to enhance he vertical interlayer hopping $t_{\perp}$ between the $d_{3z^{2}-r^{2}}$ orbitals while the nearest-neighbor intralayer hoppings $t_{3z^{2}-r^{2}}$, $t_{x^{2}-y^{2}}$, and $t_{x^{2}-y^{2}{\text{-}}3z^{2}-r^{2}}$ are somewhat suppressed as shown in Table 1.
@@ III.3 FLEX analysis on superconductivity in Sr3Ni2O5Cl2
In Fig. 4, we plot the gap function in the band representation for Sr3Ni2O5Cl2 ($\Delta n=0.01$) and La3Ni2O7 ($\Delta n=0$). For Sr3Ni2O5Cl2, the gap function of the $d_{3z^{2}-r^{2}}$ portion of the band is large with a basically $s\pm$-wave structure, i.e., -- sign in the bonding and $+$ sign in the antibonding band, as in La3Ni2O7. The gap function of the $d_{x^{2}-y^{2}}$ portion is smaller compared to La3Ni2O7 due to the smaller hybridization with $d_{3z^{2}-r^{2}}$.
@@
Figure 4: (a) The superconducting gap function for Sr3Ni2O5Cl2 ($\Delta n=0.01$) at $T=0.01$ eV in the band representation, compared with (b) that for the four-orbital model of La3Ni2O7 ($\Delta n=0$). The strength of the Wannier orbital characters is shown with the radius of the color coded circles, where the weight of the $d_{3z^{2}-r^{2}}$ orbital is indicated by either green or blue, depending on whether the band energy is above or below $E=0.1$ eV.

### v2

@@ I Introduction
Discovery of superconductivity with a maximum $T_{c}$ of above 80 K in a bilayer Ruddlesden-Popper nickelate La3Ni2O7 (Fig. 1(a)) under pressure has sparked enormous interest in the field of condensed matter physics [1]. Although under pressure, $T_{c}$ exceeding the liquid nitrogen boiling temperature, and also the origin of superconductivity likely being unconventional [2, 3, 4], have led to a huge amount of experimental and theoretical studies after the discovery [1, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 3, 4, 69, 2, 70, 71, 72, 73, 74, 75, 76, 77]. In fact, the possibility of superconductivity in this material was discussed theoretically even before its experimental discovery by some of the previous authors [78], but one viewpoint that was missing at that time was the precise crystal structure, namely, although a tetragonal symmetry was simply assumed in Ref. 78, the actual structural symmetry is orthorhombic at ambient pressure. There are now accumulating studies suggesting that tetragonal symmetry of the crystal structure may be important for the occurrence of superconductivity in La3Ni2O7 [32, 42, 41], which is accomplished by applying pressure. One speculation regarding the reason for this can be that orthorhombicity favors density waves, which suppresses superconductivity. Interestingly, also for the trilayer compound La4Ni3O10, the superconducting phase appears by applying a sufficient pressure and the tetragonal structure is realized [31, 38, 39, 40, 48]. Then, a natural motivation arises for searching materials that have tetragonal symmetry at ambient pressure, and also have electronic properties similar to that of La3Ni2O7 at high pressures.
@@
In this study, we theoretically investigate the electronic structure of Sr3Ni2O5Cl2 and its structural stability. Our phonon calculation shows that this compound with the $I4/mmm$ tetragonal structure is dynamically stable, i.e., does not have imaginary-frequency phonon modes, even at ambient pressure. The Ni-$d_{3z^{2}-r^{2}}$ orbital energy is pushed down by the crystal field where half of the apical O2- ions are replaced with a Cl- ion. By this effect, the Ni-$d_{3z^{2}-r^{2}}$ orbital becomes rather closer to the half-filling in Sr3Ni2O5Cl2 than La3Ni2O7 whereas the stoichiometric Ni-$3d$ band filling are decreased in Sr3Ni2O5Cl2 ($d^{7}$) compared with La3Ni2O7 ($d^{7.5}$). We also investigate superconductivity in this compound with the multi-orbital Hubbard model constructed from our first-principles calculation, using the fluctuation exchange (FLEX) approximation combined with the linearized Eliashberg equation. While orbital hybridization between the $t_{2g}$ and $e_{g}$ orbitals, which is enhanced in this characteristic crystal field, is detrimental for superconductivity, a nearly half-filled Ni-$d_{3z^{2}-r^{2}}$ orbital enhances superconductivity. In total, we find that Sr3Ni2O5Cl2 is a promising new candidate of bilayer-nickelate superconductors, which can possess even higher $T_{c}$ than pressurized La3Ni2O7, at ambient pressure. We also check the formation enthalpy for synthesis of this compound. Our calculation shows that the high-pressure synthesis can be a good way to experimentally produce Sr3Ni2O5Cl2.
Figure 1: (a) Crystal structure of La3Ni2O7 at high pressure. (b) Crystal structure and (c) phonon band dispersion for Sr3Ni2O5Cl2 without applying pressure. Panels (b) and (c) are shown for the space group of $I4/mmm$. Crystal structures in this paper were depicted using the VESTA software [82]. Phonon dispersions in this paper were depicted in the Brillouin zone of the conventional tetragonal lattice.
@@ II Methods
For first-principles calculation, we used the PBEsol exchange-correlation functional [83] and the projector augmented wave method as implemented in Vienna ab initio Simulation Package (VASP) [84, 85, 86, 87, 88]. Core-electron states in PAW potentials were [Kr]$4d^{10}$, [Ar]$3d^{10}$, [Ar], [Ne], and [He] for La, Sr, Ni, Cl, and O, respectively. We used a plane-wave cutoff energy of 600 eV for Kohn-Sham orbitals without including the spin-orbit coupling for simplicity. A $12\times 12\times 12$ ${\bm{k}}$-mesh was used for self-consistent-field calculations.
We performed structural optimization until the Hellmann-Feynman force becomes less than 0.01 eV Å-1 for each atom. To check the stability of the tetragonal ($I4/mmm$) phase, we calculated the phonon dispersion for several materials. For this purpose, we used the finite displacement method as implemented in the Phonopy [89, 90] software in combination with VASP. We used a $4\times 4\times 1$ ${\bm{q}}$-mesh for a conventional tetragonal unit cell. For a $4\times 4\times 1$ supercell used for finite-displacement calculations, we used a $3\times 3\times 2$ ${\bm{k}}$-mesh.
@@ III.1 Electronic and phonon band structures for Sr3Ni2O5Cl2
As described in the Introduction, a key idea in this study is to employ a mixed-anion strategy [103] for stabilizing the tetragonal $I4/mmm$ phase without applying the pressure. In fact, for Sr3Ni2O5Cl2, while the NiO2 plane is somewhat buckled; the equatorial Ni-O-Ni angle is 162 as shown in Fig. 1(b), we found that the tetragonal structure ($a=3.779$, $c=24.532$) is dynamically stable as shown by our phonon calculation (Fig. 1(c)) where no imaginary-frequency modes appear without applying the pressure. This is in sharp contrast to the case of La3Ni2O7, where the tetragonal structure is unstable at ambient pressure and high pressure is required to stabilize it [32, 42, 41]. Our phonon calculations for La3Ni2O7 also support this feature as shown in Fig. 6, where the imaginary-frequency modes appear for 0 GPa while they are absent for 20 GPa. Although we also checked the stability of the tetragonal structure for Sr3Ni2O4Cl3, where all the apical oxygens were replaced with chlorines, we found that the tetragonal structure was unstable for this case as shown in Fig. 7. We consider that one important difference among these compounds is the average valence number of the nickel atom: Ni2.5+ ($d^{7.5}$) for La3Ni2O7 and Sr3Ni2O4Cl3 while Ni3+ ($d^{7}$) for Sr3Ni2O5Cl2. Here, the size of the nickel ion becomes smaller in Sr3Ni2O5Cl2, which in general suppresses the rotational and tilting instability of octahedra in Ruddlesden-Popper phase materials.
@@
It is noteworthy that the mixed-anion strategy drastically changes the crystal field of a nickel ion. Figure 2 presents the electronic band structures and partial density of states (pDOS) for La3Ni2O7 under pressure of 20 GPa and Sr3Ni2O5Cl2 without applying the pressure shown with an orbital weight of each Ni-$d$ orbital. We summarize important differences between two materials from the viewpoint of the crystal field as follows.
@@
As mentioned above, the average valence number of the nickel atom is different: Ni2.5+ ($d^{7.5}$) for La3Ni2O7 and Ni3+ ($d^{7}$) for Sr3Ni2O5Cl2. Thus, the number of $d$ electrons for Sr3Ni2O5Cl2 is less than that for La3Ni2O7. Nevertheless, the Ni-$d_{3z^{2}-r^{2}}$ orbital is rather closer to the half-filling in Sr3Ni2O5Cl2 as shown in Figs. 2(b)(h). This is because of the characteristic crystal field for the nickel atom in Sr3Ni2O5Cl2, where a half of apical O2- ions is replaced with a Cl- ion, which pushes down the orbital energy of the Ni-$d_{3z^{2}-r^{2}}$ orbital. As a result, the number of $d_{x^{2}-y^{2}}$ electrons is decreased in Sr3Ni2O5Cl2. On the other hand, Ni-$d_{xy/yz/zx}$ orbitals, which we hereafter call them the $t_{2g}$ orbitals while it is not a rigorous term for this symmetry, raise their energy levels relative to the Ni-$d_{3z^{2}-r^{2}}$ orbital. This is because the present crystal field strongly pushes down the Ni-$d_{3z^{2}-r^{2}}$ orbital level as we have stated above. Consequently, the strong DOS peak of the $t_{2g}$ orbitals lies just below the Fermi energy. In addition, the crystal field is strongly asymmetric with respect to the $xy$ plane (i.e., $+z$ and $-z$ are strongly inequivalent), which allows a sizable hybridization among $e_{g}$ ($d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$) and $t_{2g}$ orbitals. In the next section, we will discuss how these characteristics can affect superconductivity in this system.
@@ III.2 Model construction
In this and the next sections, we compared Sr3Ni2O5Cl2 without applying the pressure and La3Ni2O7 under pressure of 20 GPa, both of which have the $I4/mmm$ tetragonal structure in our calculation. We constructed a ten-orbital model consisting of five $3d$ orbitals centered at two Ni sites per unit cell. Figure 3(a) shows superposed band structures given by first-principles calculation and Wannier interpolation, where precise fitting around the Fermi level was achieved. Important parameter values extracted by Wannierization are given in Table 1, compared with those of four-orbital models consisting only of $e_{g}$ orbitals. The smaller differences in extracted parameters between the ten- and four-orbital models for La3Ni2O7 originate from the fact that the $e_{g}$ and $t_{2g}$ bands are well separated compared with Sr3Ni2O5Cl2, where the $t_{2g}$ and $e_{g}$ bands are rather entangled, as shown in Fig. 2. We found a relatively large hopping amplitude between $t_{2g}$ and $e_{g}$ orbitals: e.g., the nearest-neighbor intralayer hopping along the $x$ direction between the $d_{x^{2}-y^{2}}$ and $d_{xz}$ orbitals and that between the $d_{3z^{2}-r^{2}}$ and $d_{xz}$ orbitals are $-0.296$ and $0.199$ eV, respectively. This is due to the strongly asymmetric crystal field for a NiO5Cl octahedron as mentioned in the previous section. It is also noteworthy that the relatively strong buckling of NiO2 plane in Sr3Ni2O5Cl2 seems to enhance he vertical interlayer hopping $t_{\perp}$ between the $d_{3z^{2}-r^{2}}$ orbitals while the nearest-neighbor intralayer hoppings $t_{3z^{2}-r^{2}}$, $t_{x^{2}-y^{2}}$, and $t_{x^{2}-y^{2}{\text{-}}3z^{2}-r^{2}}$ are somewhat suppressed as shown in Table 1. For Sr3Ni2O5Cl2, we also found that the onsite energies of the $d_{xy}$ and $d_{yz/zx}$ orbitals relative to that for the $d_{3z^{2}-r^{2}}$ orbital are $-0.811$ eV and $-0.840$ eV, respectively, the size of which are comparable to $\Delta E=E_{x^{2}-y^{2}}-E_{3z^{2}-r^{2}}=0.776$ eV listed in Table 1.
@@ III.3 FLEX analysis on superconductivity in Sr3Ni2O5Cl2
In Fig. 4, we plot the gap function in the band representation for Sr3Ni2O5Cl2 ($\Delta n=0.01$) and La3Ni2O7 ($\Delta n=0$). For Sr3Ni2O5Cl2, the gap function of the $d_{3z^{2}-r^{2}}$ portion of the band is large with a basically $s\pm$-wave structure, i.e., -- sign in the bonding and $+$ sign in the antibonding band, as in La3Ni2O7. The gap function of the $d_{x^{2}-y^{2}}$ portion is smaller compared to La3Ni2O7 due to the smaller hybridization with $d_{3z^{2}-r^{2}}$. The sign change between the bonding and antibonding $d_{3z^{2}-r^{2}}$ bands corresponds to the dominant contribution of the off-diagonal $d_{3z^{2}-r^{2}}$ component in the orbital representation, as shown in Fig. 4(c), which means that the pairing takes place mainly in the interlayer channel.
@@
Here, we make a brief comment on the robustness of our results against the interaction parameters. We also calculated $\lambda$ for Sr3Ni2O5Cl2 at stoichiometry using the interaction parameters evaluated with the constrained random-phase approximation (cRPA) and obtained $\lambda=0.767$. Although this value is slightly smaller than $\lambda=0.806$ shown in Fig. 4(c), it is still sufficiently higher than $\lambda$ for pressurized La3Ni2O7 at stoichiometry. The obtained interaction parameters are shown in Appendix D. We also note that, in our previous study [8], we verified that $\lambda$ for La3Ni2O7 is similar between two sets of interaction parameters: ($U,U^{\prime},J=J^{\prime}$) = (3, 2.4, 0.3) eV and those evaluated with cRPA [11].
Figure 4: (a) The superconducting gap function for Sr3Ni2O5Cl2 ($\Delta n=0.01$) at $T=0.01$ eV in the band representation, compared with (b) that for the four-orbital model of La3Ni2O7 ($\Delta n=0$). The strength of the Wannier orbital characters is shown with the radius of the color coded circles, where the weight of the $d_{3z^{2}-r^{2}}$ orbital is indicated by either green or blue, depending on whether the band energy is above or below $E=0.1$ eV. (c) Same as (a) but in the orbital representation, where the orbital-diagonal components for all $d$ orbitals and the orbital-off-diagonal component for the $d_{3z^{2}-r^{2}}$ orbital are shown.